OpenStudy (anonymous):

Which binomial is a factor of f(x) = x4 - 3x3 - 28x2 + 36x + 144? (x - 6) (x - 4) (x - 2) (x + 3)

OpenStudy (misty1212):

oh wow what did you to to annoy your math teacher?

OpenStudy (anonymous):

lol haha I am behind and trust me algebra is not my strongest point lol

OpenStudy (misty1212):

i have an idea lets cheat

OpenStudy (misty1212):

apparently this is the same as \[(x+2) (x-3) (x+4) (x-6)\]check here http://www.wolframalpha.com/input/?i=+x4+-+3x3+-+28x2+%2B+36x+%2B+144

OpenStudy (misty1212):

pick \(x-6\) i do not know how you could do this on your own

OpenStudy (misty1212):

i mean i do, but it would take like forever

OpenStudy (anonymous):

omg thats awesome your a lifesaver right now i know so time consuming

OpenStudy (misty1212):

glad i can help you ;)

OpenStudy (anonymous):

According to the rational root theorem, which is not a possible rational root 5x3+3x2-2x+15=0 A. -3 B. 1/5 C.1/3 D.5

OpenStudy (misty1212):

the denominator needs to be a factor of the leading coefficient (\(5\)) 3 is not a factor of 5

OpenStudy (anonymous):

so Its 5 right?

OpenStudy (anonymous):

or 1/5

OpenStudy (misty1212):

no it says which is NOT the one that is NOT is \(\frac{1}{3}\)

OpenStudy (anonymous):

oh your right thanks so much

OpenStudy (misty1212):

:)

OpenStudy (anonymous):

Which is the factored form of x3-10x2+27x-18=0 A. (x+1)(x-3)(x-6)=0 B. (x+1)(x-2)(x-6)=0 C. (x-1)(x-3)(x-6)=0 D. (x+1)(x-2)(x+9)=0

OpenStudy (misty1212):

who knows

OpenStudy (misty1212):

the shadow knows http://www.wolframalpha.com/input/?i=+x3-10x2%2B27x-18%3D0

OpenStudy (anonymous):

Im not seeing it lol

OpenStudy (anonymous):

i seen it lol ohh myy

OpenStudy (anonymous):

How many complex roots does 5x6+x4-3=0 3 4 5 6